# Is Meta-training Really Necessary for Molecular Few-Shot Learning ?

### v1

@@ Abstract
Machine Learning, ICML
Figure 1: Performances of different few-shot learning methods evaluated with library screening (HTS) out-of-distribution tasks from the Therapeutic Data Commons dataset. Models are evaluated with the hitrate in the top-5% of the predictions (left). A ranking of the models on these tasks across different proportions of selected molecules reveals the competitiveness of fine-tuning baselines such as the linear probe (l-probe) or our proposed quadratic probe (q-probe) in comparisons to meta-learning methods (right). Complete results are available in the section 4.
@@ 1 Introduction
Drug discovery is a process that aims to answer the following question: “How do we design molecules with biological activity on a specific therapeutic target and other necessary properties to become a drug?”. To identify such compounds, scientists rely on a combination of experimental and computational methods (Xu & Hagler, 2002; Eckert & Bajorath, 2007). This lengthy and costly process takes on average 12 years and 1.8 billion dollars to bring a new drug to the market (Pammolli et al., 2019; Arrowsmith, 2011). To reduce the cost associated with this process and to speed up the development of new drugs, machine learning methods are often used to guide the discovery of new compounds (Gilmer et al., 2017; Chen et al., 2018). A common task for these models is to predict the activity of a compound on a given target exploiting the compound’s structural or physio-chemical properties, referred to as quantitative structure-activity relationship (QSAR) (Tropsha, 2010). QSAR models can also be used as oracles to guide generative algorithms into generating compounds while optimizing their properties (Bengio et al., 2021; Blaschke et al., 2020; Gao et al., 2022).
@@
Meta-learning. Meta-learning corresponds to a family of methods pre-trained to ’learn to learn’ from a small set of labelled examples (support set), directly conditioning their predictions on this support set (Vinyals et al., 2016).
Hence, these algorithms must pretrain a model specific to few-shot learning tasks. This implies that different pre-trained models must be maintained to perform few-shot and standard classification. Furthermore, most meta-learning algorithms use specific training procedures, as the training and test-time scenarios must be similar. Controlling the training procedure of these methods is hence necessary to use them. This can be problematic in a black-box setting, where only the predictions of the model can be queried through an API, and the model’s weights cannot be accessed (Ji et al., 2021; Ouali et al., 2023). Using these methods in a black-box scenario would require imposing the training procedure on the server, which is not always possible, or would limit the number of models that could be used. Yet this black-box setting is especially valuable in drug discovery, where data is highly sensitive, and sharing a dataset or a pre-trained model on a private dataset is not always possible.
Competitive fine-tuning methods. Fine-tuning methods correspond to using a model pre-trained in a standard supervised setting and fine-tuning it for few-shot classification. Furthermore, different backbones, pre-trained with different pre-training objectives, can be adapted with such methods. Hence, a wider range of models can be used. In other domains, such as computer vision, meta-learning methods were shown to obtain only marginal improvements over more straightforward fine-tuning methods, using models pre-trained in a standard supervised setting (Chen et al., 2019; Luo et al., 2023). Some of these results were made possible thanks to the availability of foundational models pre-trained on large datasets available in these fields (Radford et al., 2021; Brown et al., 2020). However, developing such foundational models is a non-trivial task in the context of drug discovery, and is still an active research field (Seidl et al., 2023; Ahmad et al., 2022). In drug discovery, it was observed that fine-tuning approaches yielded subpar performances (Stanley et al., 2021), a finding seemingly at odds with existing results in vision applications.
Contributions
We re-visit fine-tuning black-box models using a simple Multitask backbone to perform QSAR few-shot learning. Our contributions could be summarized as follows:
@@ 2 Related Work
Notations. Assume that we are given a base dataset $\mathcal{D}_{\text{base}}=\{\mathcal{D}_{\text{base}}^{t}\}_{t\in\mathbf{T}_{\text{base}}}$, where $\mathbf{T}_{\text{base}}$ is a set of tasks. For a task $t$, we have $\mathcal{D}_{\text{base}}^{t}=\{{\mathbf{x}}_{i},y_{i}\}_{0\leq i\leq N^{t}_{\text{base}}}$, where ${\mathbf{x}}_{i}$ is a molecule and $y_{i}\in\{0,1\}$ denotes its label.
We then have access to a test dataset defined similarly: $\mathcal{D}_{\text{test}}=\{\mathcal{D}_{\text{test}}^{t}\}_{t\in\mathbf{T}_{\text{novel}}}$ where $\mathbf{T}_{\text{novel}}$ is the set of new tasks ($\mathbf{T}_{\text{novel}}\cap\mathbf{T}_{\text{base}}=\emptyset$). Finally, for a task $t$, $\mathcal{D}_{\text{test}}^{t}$ is divided into a set of labelled data that we will use to adapt the model, i.e., support set $\mathcal{S}_{t}=\{{\mathbf{x}}_{i},y_{i}\}_{0\leq i\leq N_{\text{support}}}$, and an unlabelled set whose labels are to be predicted, i.e., query set $\mathcal{Q}_{t}=\{{\mathbf{x}}_{i},y_{i}\}_{0\leq i\leq N_{\text{query}}}$.
@@
The work of Stanley et al. (2021) introduced a benchmark for molecular few-shot learning, FS-mol, and evaluated various methods. For instance, the authors trained an adaptation of Protonet (Snell et al., 2017) with a graph neural network (GNN). Besides, the authors pre-trained a graph neural network (GNN) with the MAML algorithm (Finn et al., 2017), a model aiming at learning a good initialization of the model’s weights, such that the model can be quickly adapted to new tasks. Building on the MAML paradigm, the property-aware relation network (PAR) (Wang et al., 2021) builds a similarity graph between the molecules of the support set and the queried molecule and extracts from this graph the molecule’s predicted class. In recent work, Chen et al. (2023) proposed a method separating the parameters $\theta_{meta}$ that are updated during meta-training, while the other parameters $\theta_{\textrm{adapt}}$ are only updated during the adaptation. The back-propagation on each task is done by using Cauchy’s implicit function theorem, which requires estimating the hessian of the loss $\frac{\partial^{2}\mathcal{L}_{\textrm{support}}}{\partial\theta_{\textrm{adapt}}^{2}}$ on the support set. Finally, some work evaluated the effect of leveraging a large set of context molecules to construct similarity measures between the support set’s molecules and the queried molecule using modern hopfield networks (Schimunek et al., 2023).
@@
Fine-tuning methods. Among the baselines presented by Stanley et al. (Stanley et al., 2021), some of them used a backbone trained without relying on meta-training. First, the authors fine-tuned a GNN trained with a multitask objective on the support set of the novel task, consisting of a GNN common trunk and a 2-layer perceptron task-specific head. At test time, both the common-trunk and the task-specific parameters are fine-tuned on the support set. A second approach used the molecular attention transformer (MAT) (Maziarka et al., 2019) and fine-tuned it on the support set of the novel task. Both methods achieved performances on FS-mol below the aforementioned meta-learning methods. Furthermore, they need to fine-tune the whole model, which can be computationally expensive and requires the user to access the model’s weights at test time. Hence, these methods cannot be directly used in a black-box setting. Finally, the work of Seidl et al. (2023) proposed a multi-modal model leveraging both the molecular graph and a textual representation of the assays to make predictions, enabling the creation of zero-shot models. The authors evaluated the model on FS-mol with a linear probe. However, to train their model, the support set of all tasks was integrated into the pre-training set, which breaks the supposition $\mathbf{T}_{novel}\cap\mathbf{T}_{base}=\emptyset$. This evaluation corresponds to an assessment of the model’s ability to generalize on previously seen tasks with few labels, not the model’s ability to adapt to unseen tasks.
@@ 3.1 Model pre-training
Various methods exist to pre-train models for drug discovery (Hu* et al., 2020), which can usually be divided into supervised and self-supervised methods. Supervised pre-training refers to methods training a model on a dataset of molecules relying on the association between the input and labels. In contrast, self-supervised pre-training refers to methods training a model on a dataset of molecules without using or knowing the labels. While self-supervised pre-training enables the use of larger data sets, supervised pre-training tends to be more effective when the labels used during training correspond to assays similar to those of the downstream task (Sun et al., 2022). In this work, we pursue the supervised setting via a multitask model based on the training set of the FS-mol benchmark (Stanley et al., 2021). We focus our work on QSAR modelling, evaluating the ability of the pre-training backbone to adapt to new tasks.
The GNN backbone $f_{\theta}$ will take the molecular graph and the fingerprints as input, based on a principal neighbourhood aggregation model (Corso et al., 2020). We add a task-specific head $g_{\phi}$ to the backbone, a one-layer binary classifier during the pre-training. Once the training is finished, we only keep the parameters $\theta$ of the backbone. The molecular embeddings passed to the few-shot classifier are obtained by
${\mathbf{z}}_{i}=\frac{f_{\theta}({\mathbf{x}}_{i})}{\lVert f_{\theta}({\mathbf{x}}_{i})\rVert}=\frac{f_{\theta}({\mathcal{G}}_{i},{\mathbf{f}}_{i})}{\lVert f_{\theta}({\mathcal{G}}_{i},{\mathbf{f}}_{i})\rVert},$
with ${\mathcal{G}}_{i}$ denoting the molecular graph, ${\mathbf{f}}_{i}$ the molecular fingerprint.
@@ 3.2 Multitask Linear Probing
As explained in the section 2, multitask models were shown to be outperformed by most other methods (Stanley et al., 2021). Still, we show that we can obtain better classification capabilities by following the paradigm of linear probing (Chen et al., 2019). Linear probing corresponds to a model training the parameters ${\mathbf{w}}_{k},\lVert{\mathbf{w}}_{k}\rVert=1$ for each class $k\in C=\{0,\ldots,n_{\textrm{class}}\}$. The predicted probability of a molecule ${\mathbf{x}}_{i}$ to belong to the class $k\in C$ is then given by:
$p_{i,k}=\frac{\exp{\left(\tau\langle{\mathbf{z}}_{i},{\mathbf{w}}_{k}\rangle+b_{k}\right)}}{\sum_{k^{\prime}\in C}\exp{\left(\tau\langle{\mathbf{z}}_{i},{\mathbf{w}}_{k^{\prime}}\rangle+b_{k^{\prime}}\right)}}$ (1)
where $b_{k}$ is the model’s bias, $\tau$ is a temperature hyperparameter, and for the clarity of the notation, we omitted the dependency on the task ${\mathbf{t}}$. The linear probe’s parameters are then optimized to minimize the cross-entropy loss on the support set.
@@ 3.3 Quadratic probing
Motivation We propose an extension of the aforementioned linear probe, which is more competitive as the support set grows. Instead of the cosine similarity used in the linear probe, we propose to use the Mahalanobis distance between the query points and the class prototypes:
$\lVert{\mathbf{z}}_{i}-{\mathbf{w}}_{k}\rVert^{2}_{\Sigma_{k}^{-1}}=\left({\mathbf{z}}_{i}-{\mathbf{w}}_{k}\right)^{T}\Sigma_{k}^{-1}\left({\mathbf{z}}_{i}-{\mathbf{w}}_{k}\right),$ (2)
where $\Sigma_{k}$ is a positive semi-definite (PSD) matrix. For simplicity, we will refer to ${\mathbf{M}}_{k}\triangleq\Sigma_{k}^{-1}$ as the precision matrix of the distribution to simplify the notations. The predicted probability of a molecule ${\mathbf{x}}_{i}$ to belong to the class $k\in C$ can be then obtained as:
$\centering p_{i,k}\triangleq\frac{\exp{\left(-\lVert{\mathbf{z}}_{i}-{\mathbf{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}\right)}}{\sum_{k^{\prime}\in C}\exp{\left(-\lVert{\mathbf{z}}_{i}-{\mathbf{w}}_{k^{\prime}}\rVert^{2}_{{\mathbf{M}}_{k}}\right)}}.\@add@centering$ (3)
The Mahalanobis distance measures the distance between a point and a normal distribution of mean ${\mathbf{w}}_{k}$ and covariance matrix $\Sigma_{k}$. This distance is more expressive than the cosine similarity, as it considers the covariance matrix of the class. In fact, one can see it as generalization of the linear probe, which corresponds to choosing a covariance matrix proportional to identity. Figure 2 depicts an illustration of the difference between linear probe and our quadratic probe.
Figure 2: Illustration of the difference between using the linear probe approach (left) and classifiers based on a Mahalanobis distance (right).
In the next section, we show that directly minimizing the cross-entropy loss over ${\mathbf{M}}_{k}$ yields degenerate solutions, in which the Frobenius norms of the covariance matrices go to infinity. To mitigate this issue, we propose a surrogate objective, which corresponds approximately to augmenting the cross-entropy loss with a regularization term based on the Frobinous norms of the covariance matrices, while yielding closed-form solutions at each optimization step w.r.t ${\mathbf{M}}_{k}$. We proceed with block-coordinate descent approach to optimize our regularized quadratic-probe loss, updating the parameters $\{{\mathbf{w}}_{k}\}_{k\in\{0,1\}}$ and $\{{\mathbf{M}}_{k}\}_{k\in\{0,1\}}$ separately.
@@
Let $\Theta=\{{\mathbf{w}}_{k},{\mathbf{M}}_{k}\}_{k\in\{0,1\}}$, and we will note $p_{\Theta}(k|{\mathbf{z}})=p_{i,k}$ as described in Equation 3 (to highlight the dependency to the parameters $\Theta$).
If the samples from both classes are linearly separable, we can construct a set of parameters, $\Theta(\lambda)=\{{\mathbf{w}}_{k},{\mathbf{M}}_{k}(\lambda)\}_{k\in\{0,1\}}$ such that:
$\begin{cases}\forall i\in\{0,\ldots,N\},\quad\lim_{\lambda\to+\infty}\mathcal{L}_{ce}({\mathbf{z}}_{i},y_{i},{\Theta(\lambda)})=0,\\ \forall k\in\{0,1\},\quad\lim_{\lambda\to+\infty}\lVert{\mathbf{M}}_{k}(\lambda)\rVert_{F}=+\infty,\end{cases}$
where $\mathcal{L}_{ce}({\mathbf{z}}_{i},y_{i},\Theta)=-\log p_{\Theta}(y_{i}|{\mathbf{z}}_{i})$ is the point-wise cross-entropy loss w.r.t our model.
The proof of proposition 3.1 is relegated to Appendix A.0.1. This proposition relies on the assumption that the data points are linearly separable. In the few-shot learning setting, this assumption is often verified. Indeed, as the data dimension is much larger than the number of data points, Cover’s theorem (Cover, 1965) states that the points are almost certainly linearly separable. Another argument supporting the assumption of linear separability comes from an approximation often made in high-dimensional spaces:
$\text{if }d\gg N,\quad\forall i,j\in\{0,\ldots,N\}^{2}\quad z_{i}^{t}z_{j}=0.$
Then, the vector $v=\sum_{i=0}^{N}y_{i}z_{i}-(1-y_{i})z_{i}$ defines a hyperplane separating data points in both classes.
Block coordinate optimization and modified loss.
Instead of optimizing the parameters of the quadratic probe by gradient descent, we optimize the parameters of the model with a block coordinate descent, updating the parameters $\{{\mathbf{w}}_{k}\}_{k\in\{0,1\}}$ and $\{{\mathbf{M}}_{k}\}_{k\in\{0,1\}}$ separately. In particular, we will perform one gradient descent step on $\{{\mathbf{w}}_{k}\}_{k\in\{0,1\}}$ using the cross-entropy loss and then minimize a modified loss w.r.t $\{{\mathbf{M}}_{k}\}_{k\in\{0,1\}}$, that does not suffer from the issue described in proposition 3.1. In the following section, $\{{\mathbf{w}}_{k}\}_{k\in\{0,1\}}$ is considered fixed.
@@
$\mathcal{L}_{ce}(\mathcal{S},\Theta)=\frac{1}{|\mathcal{S}|}\sum_{i\in\mathcal{S}}\left(f_{1}\left({\mathbf{z}}_{i},y_{i},\Theta\right)+f_{2}\left({\mathbf{z}}_{i},\Theta\right)\right),$ (4)
@@
$f_{1}\left({\mathbf{z}}_{i},y_{i},\Theta\right)\triangleq\lVert{\mathbf{z}}-{\mathbf{w}}_{y_{i}}\rVert^{2}_{{\mathbf{M}}_{y_{i}}},$ (5)
$f_{2}\left({\mathbf{z}}_{i},\Theta\right)\triangleq\log\left(\sum_{k\in\{0,1\}}\exp\left(-\lVert{\mathbf{z}}-{\mathbf{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}\right)\right).$
@@
For all $i\in\mathcal{S}$, $f_{1}\left({\mathbf{z}}_{i},y_{i},\Theta\right)$ is minimized when $\lVert{\mathbf{M}}_{k}\rVert_{F}\to 0$, and $f_{2}\left({\mathbf{z}}_{i},y_{i},\Theta\right)$ as $\lVert{\mathbf{M}}_{k}\rVert_{F}\to+\infty$.
The function $f_{2}$ is hence responsible for the degenerate solutions obtained while minimizing the cross-entropy loss. We propose to replace $f_{2}$ with a different function $\tilde{f}_{2}$, playing the same role pushing $\lVert{\mathbf{M}}_{k}\rVert_{F}$ towards infinity, but with a less aggressive behaviour. The new loss obtained $f_{1}+\tilde{f}_{2}$ admits a closed-form solution on ${\mathbf{M}}_{k}$, for a fixed ${\mathbf{w}}_{k}$, which corresponds to the estimator of the empirical covariance matrix’s inverse. The expression of $\tilde{f}_{2}$ is then:
$\tilde{f}_{2}\left(\Theta\right)\triangleq-\sum_{k\in\{0,1\}}\frac{\lvert\mathcal{S}_{k}\rvert\log\det{({\mathbf{M}}_{k})}}{\lvert\mathcal{S}\rvert}.$

### v2

# A Strong Baseline for Molecular Few-Shot Learning

@@ 1 Introduction
Drug discovery is a process that aims to answer the following question: ”How do we design molecules with biological activity on a specific therapeutic target and other necessary properties to become a drug?”. To identify such compounds, scientists rely on a combination of experimental and computational methods (Xu & Hagler, 2002; Eckert & Bajorath, 2007). This lengthy and costly process takes on average 12 years and 1.8 billion dollars to bring a new drug to the market (Pammolli et al., 2019; Arrowsmith, 2011). To reduce the cost associated with this process and to speed up the development of new drugs, machine learning methods are often used to guide the discovery of new compounds (Gilmer et al., 2017; Chen et al., 2018). A common task for these models is to predict the activity of a compound on a given target exploiting the compound’s structural or physio-chemical properties, referred to as quantitative structure-activity relationship (QSAR) (Tropsha, 2010). QSAR models can also be used as oracles to guide generative algorithms into generating compounds while optimizing their properties (Bengio et al., 2021; Blaschke et al., 2020; Gao et al., 2022).
@@
Meta-learning. Meta-learning corresponds to a family of methods pre-trained to ’learn to learn’ from a small set of labelled examples (support set), directly conditioning their predictions on this support set (Vinyals et al., 2016). Hence, these algorithms must pretrain a model specific to few-shot learning tasks. This implies that different pre-trained models must be maintained to perform few-shot and standard classification. Furthermore, most meta-learning algorithms use specific training procedures, as the training and test-time scenarios must be similar. Controlling the training procedure of these methods is hence necessary to use them. This can be problematic in a black-box setting, where only the predictions of the model can be queried through an API, and the model’s weights cannot be accessed (Ji et al., 2021; Ouali et al., 2023). Using these methods in a black-box scenario would require imposing the training procedure on the server, which is not always possible, or would limit the number of models that could be used. Yet this black-box setting is especially valuable in drug discovery, where data is highly sensitive, and sharing a dataset or a pre-trained model on a private dataset is not always possible (Heyndrickx et al., 2024; Pejo et al., 2022; Smajić et al., 2023).
Competitive fine-tuning methods. Fine-tuning methods involve using a model pre-trained in a standard supervised setting and then fine-tuning it for few-shot classification. Unlike meta-learning, fine-tuning approaches utilize pre-trained models that can be applied to other tasks in standard machine learning, beyond few-shot classification, hence, a wider range of models can be used. In other domains, such as computer vision, meta-learning methods were shown to obtain only marginal improvements over more straightforward fine-tuning methods, using models pre-trained in a standard supervised setting (Chen et al., 2019; Luo et al., 2023). Some of these results were made possible thanks to the availability of foundational models pre-trained on large datasets available in these fields (Radford et al., 2021; Brown et al., 2020). However, developing such foundational models is a non-trivial task in the context of drug discovery, and is still an active research field (Seidl et al., 2023; Ahmad et al., 2022; Méndez-Lucio et al., 2022). In drug discovery, it was observed that fine-tuning approaches yielded subpar performances (Stanley et al., 2021), a finding seemingly at odds with existing results in vision applications.
Contributions. We re-visit fine-tuning black-box models using a simple Multitask backbone to perform QSAR few-shot learning. Our contributions could be summarized as follows:
@@ 2 Related Work
Notations. Assume that we are given a base dataset $\mathbb{D}_{\text{base}}=\{\mathbb{D}_{\text{base}}^{t}\}_{t\in{\mathbb{T}}_{\text{base}}}$, where ${\mathbb{T}}_{\text{base}}$ is a set of tasks. For a task $t$, we have $\mathbb{D}_{\text{base}}^{t}=\{{\bm{x}}_{i},y_{i}\}_{0\leq i\leq N^{t}_{\text{base}}}$, where ${\bm{x}}_{i}$ is a molecule and $y_{i}\in\{0,1\}$ denotes its label.
We then have access to a test dataset defined similarly: $\mathbb{D}_{\text{test}}=\{\mathbb{D}_{\text{test}}^{t}\}_{t\in{\mathbb{T}}_{\text{novel}}}$ where ${\mathbb{T}}_{\text{novel}}$ is the set of new tasks (${\mathbb{T}}_{\text{novel}}\cap{\mathbb{T}}_{\text{base}}=\emptyset$). Finally, for a task $t$, $\mathbb{D}_{\text{test}}^{t}$ is divided into two sets: $\mathcal{S}_{t},{\mathcal{Q}}_{t}$ ($\mathcal{S}_{t}\cap{\mathcal{Q}}_{t}=\emptyset$) defining a set of labelled data that we will use to adapt the model, i.e., support set $\{{\bm{x}}_{i},y_{i}\}_{i\in\mathcal{S}_{t}}$, and an unlabelled set whose labels are to be predicted, i.e., query set $\{{\bm{x}}_{i},y_{i}\}_{i\in{\mathcal{Q}}_{t}}$.
@@
The work of Stanley Stanley et al. (2021) introduced a benchmark for molecular few-shot learning, FS-mol, and evaluated various methods. For instance, the authors trained an adaptation of Protonet (Snell et al., 2017) with a graph neural network (GNN). Besides, the authors pre-trained a graph neural network (GNN) with the MAML algorithm (Finn et al., 2017), a model aiming at learning a good initialization of the model’s weights, such that the model can be quickly adapted to new tasks. Building on the MAML paradigm, the property-aware relation network (PAR) (Wang et al., 2021) builds a similarity graph between the molecules of the support set and the queried molecule and extracts from this graph the molecule’s predicted class. In recent work, Chen et al. (2023) proposed a method separating the parameters $\theta_{meta}$ that are updated during meta-training, while the other parameters $\theta_{\textrm{adapt}}$ are only updated during the adaptation. The back-propagation on each task is done by using Cauchy’s implicit function theorem, which requires estimating the hessian of the loss $\frac{\partial^{2}\mathcal{L}_{\textrm{support}}}{\partial\theta_{\textrm{adapt}}^{2}}$ on the support set. Finally, some work evaluated the effect of leveraging a large set of context molecules to construct similarity measures between the support set’s molecules and the queried molecule using modern hopfield networks (Schimunek et al., 2023).
@@
Fine-tuning methods. Among the baselines presented by Stanley et al. (Stanley et al., 2021), some of them used a backbone trained without relying on meta-training. First, the authors fine-tuned a GNN trained with a multitask objective. This model consists in a GNN common trunk followed by a 2-layer task-specific head. At test time a new task-specific head is added corresponding to the new task to adapt on. Both the common-trunk and the task-specific parameters are fine-tuned on the support set. A second approach used the molecular attention transformer (MAT) (Maziarka et al., 2019) and fine-tuned it on the support set of the novel task. Both methods achieved performances on FS-mol below the aforementioned meta-learning methods. Furthermore, they need to fine-tune the whole model, which can be computationally expensive and requires the user to access the model’s weights at test time. Hence, these methods cannot be directly used in a black-box setting. In the recent years, a significant effort was made to build foundational models able to handle molecular data (Méndez-Lucio et al., 2022; Beaini et al., 2024; Khan et al., 2024; Xia et al., 2023), but very few proposed methods assess the ability of these models to perform few-shot classification. Notably, the work of Seidl et al. (2023) proposed a multi-modal model leveraging both the molecular graph and a textual representation of the assays to make predictions, enabling the creation of zero-shot models. The authors evaluated the model on FS-mol with a linear probe. However, to train their model, the support set of all tasks was integrated into the pre-training set, which breaks the assumption ${\mathbb{T}}_{novel}\cap{\mathbb{T}}_{base}=\emptyset$. This evaluation corresponds to an assessment of the model’s ability to generalize on previously seen tasks with few labels, not the model’s ability to adapt to unseen tasks.
@@ 3.1 Model pre-training
Various methods exist to pre-train models for drug discovery (Hu* et al., 2020), which can usually be divided into supervised and self-supervised methods. Supervised pre-training refers to methods training a model on a dataset of molecules relying on the association between the input and labels. In contrast, self-supervised pre-training refers to methods training a model on a dataset of molecules without using or knowing the labels. While self-supervised pre-training enables the use of larger data sets, supervised pre-training tends to be more effective when the labels used during training correspond to assays similar to those of the downstream task (Sun et al., 2022). In this work, we pursue the supervised setting via a multitask model based on the training set of the FS-mol benchmark (Stanley et al., 2021). We focus our work on QSAR modeling, evaluating the ability of the pre-training backbone to adapt to new tasks. An evaluation of pretrained models originating from the molecular representation learning literature is available in Appendix A.6
The GNN backbone $f_{\theta}$ will take the molecular graph and the fingerprints as input, based on a principal neighborhood aggregation model (Corso et al., 2020). We add a task-specific head $g_{\phi}$ to the backbone, a one-layer binary classifier during the pre-training. Once the training is finished, we only keep the parameters $\theta$ of the backbone. The molecular embeddings passed to the few-shot classifier are obtained by
${\bm{z}}_{i}=\frac{f_{\theta}({\bm{x}}_{i})}{\lVert f_{\theta}({\bm{x}}_{i})\rVert}=\frac{f_{\theta}({\mathcal{G}}_{i},{\bm{f}}_{i})}{\lVert f_{\theta}({\mathcal{G}}_{i},{\bm{f}}_{i})\rVert},$
with ${\mathcal{G}}_{i}$ denoting the molecular graph, ${\bm{f}}_{i}$ the molecular fingerprint.
@@ 3.2 Multitask Linear Probing
As explained in the section 2, multitask models were shown to be outperformed by most other methods (Stanley et al., 2021). Still, we show that we can obtain better classification capabilities by following the paradigm of linear probing (Chen et al., 2019). Linear probing corresponds to a model training the parameters ${\bm{w}}_{k},\lVert{\bm{w}}_{k}\rVert=1$ for each class $k\in C=\{0,\ldots,n_{\textrm{class}}\}$. The predicted probability of a molecule ${\bm{x}}_{i}$ to belong to the class $k\in C$ is then given by:
$p_{i,k}=\frac{\exp{\left(\tau\langle{\bm{z}}_{i},{\bm{w}}_{k}\rangle+b_{k}\right)}}{\sum_{k^{\prime}\in C}\exp{\left(\tau\langle{\bm{z}}_{i},{\bm{w}}_{k^{\prime}}\rangle+b_{k^{\prime}}\right)}}$ (1)
where $b_{k}$ is the model’s bias, $\tau$ is a temperature hyperparameter, and for the clarity of the notation, we omitted the dependency on the task ${\bm{t}}$. The linear probe’s parameters are then optimized to minimize the cross-entropy loss on the support set.
@@ 3.3 Quadratic probing
Figure 1: Values of the logits produced by two different classifiers: a linear probe (left) and a quadratic probe based on the Mahalanobis distance (right), in a two-dimensional feature space with two classes.
Motivation. We propose an extension of the aforementioned linear probe, which is more competitive as the support set grows. Instead of the cosine similarity used in the linear probe, we propose to use the Mahalanobis distance between the query points and the class prototypes:
$\lVert{\bm{z}}_{i}-{\bm{w}}_{k}\rVert^{2}_{\Sigma_{k}^{-1}}=\left({\bm{z}}_{i}-{\bm{w}}_{k}\right)^{T}\Sigma_{k}^{-1}\left({\bm{z}}_{i}-{\bm{w}}_{k}\right),$ (2)
where $\Sigma_{k}$ is a positive semi-definite (PSD) matrix. For simplicity, we will refer to ${\mathbf{M}}_{k}\triangleq\Sigma_{k}^{-1}$ as the precision matrix of the distribution to simplify the notations.
The predicted probability of a molecule ${\bm{x}}_{i}$ to belong to the class $k\in C$ can be then obtained as:
$\centering p_{i,k}\triangleq\frac{\exp{\left(-\lVert{\bm{z}}_{i}-{\bm{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}\right)}}{\sum_{k^{\prime}\in C}\exp{\left(-\lVert{\bm{z}}_{i}-{\bm{w}}_{k^{\prime}}\rVert^{2}_{{\mathbf{M}}_{k}}\right)}}.\@add@centering$ (3)
The Mahalanobis distance measures the distance between a point and a normal distribution of mean ${\bm{w}}_{k}$ and covariance matrix $\Sigma_{k}$. This distance is more expressive than the cosine similarity, as it considers the covariance matrix of the class. In fact, one can see it as generalization of the linear probe, which corresponds to choosing a covariance matrix proportional to identity. Figure 1 depicts an illustration of the difference between linear probe and our quadratic probe.
In the next section, we show that directly minimizing the cross-entropy loss over ${\mathbf{M}}_{k}$ yields degenerate solutions, in which the Frobenius norms of the covariance matrices go to infinity. To mitigate this issue, we propose a surrogate objective, which corresponds approximately to augmenting the cross-entropy loss with a regularization term based on the Frobinous norms of the covariance matrices, while yielding closed-form solutions at each optimization step w.r.t ${\mathbf{M}}_{k}$. We proceed with block-coordinate descent approach to optimize our regularized quadratic-probe loss, updating the parameters $\{{\bm{w}}_{k}\}_{k\in\{0,1\}}$ and $\{{\mathbf{M}}_{k}\}_{k\in\{0,1\}}$ separately.
@@
Let $\Theta=\{{\bm{w}}_{k},{\mathbf{M}}_{k}\}_{k\in\{0,1\}}$, and we will note $p_{\Theta}(k|{\bm{z}})=p_{i,k}$ as described in Equation 3 (to highlight the dependency to the parameters $\Theta$).
If the samples from both classes are linearly separable, we can construct a set of parameters, $\Theta(\lambda)=\{{\bm{w}}_{k},{\mathbf{M}}_{k}(\lambda)\}_{k\in\{0,1\}}$ such that:
$\begin{cases}\forall i\in\{0,\ldots,N\},\quad\lim_{\lambda\to+\infty}\mathcal{L}_{ce}({\bm{z}}_{i},y_{i},{\Theta(\lambda)})=0,\\ \forall k\in\{0,1\},\quad\lim_{\lambda\to+\infty}\lVert{\mathbf{M}}_{k}(\lambda)\rVert_{F}=+\infty,\end{cases}$
where $\mathcal{L}_{ce}({\bm{z}}_{i},y_{i},\Theta)=-\log p_{\Theta}(y_{i}|{\bm{z}}_{i})$ is the point-wise cross-entropy loss w.r.t our model.
The proof of proposition 3.1 is relegated to Appendix A.2. This proposition relies on the assumption that the data points are linearly separable. In the few-shot learning setting, this assumption is often verified. Indeed, as the data dimension is much larger than the number of data points, Cover’s theorem (Cover, 1965) states that the points are almost certainly linearly separable. Another argument supporting the assumption of linear separability comes from an approximation often made in high-dimensional spaces: $\text{if }d\gg N,\quad\forall i,j\in\{0,\ldots,N\}^{2}\quad{\bm{z}}_{i}^{t}{\bm{z}}_{j}=0.$ Then, the vector ${\bm{v}}=\sum_{i=0}^{N}y_{i}{\bm{z}}_{i}-(1-y_{i}){\bm{z}}_{i}$ defines a hyperplane separating data points in both classes.
Block coordinate optimization and modified loss. Instead of optimizing the parameters of the quadratic probe by gradient descent, we optimize the parameters of the model with a block coordinate descent, updating the parameters $\{{\bm{w}}_{k}\}_{k\in\{0,1\}}$ and $\{{\mathbf{M}}_{k}\}_{k\in\{0,1\}}$ separately. In particular, we will perform one gradient descent step on $\{{\bm{w}}_{k}\}_{k\in\{0,1\}}$ using the cross-entropy loss and then minimize a modified loss w.r.t $\{{\mathbf{M}}_{k}\}_{k\in\{0,1\}}$, that does not suffer from the issue described in proposition 3.1. In the following section, $\{{\bm{w}}_{k}\}_{k\in\{0,1\}}$ is considered fixed.
@@
$\mathcal{L}_{ce}(\mathcal{S},\Theta)=\frac{1}{|\mathcal{S}|}\sum_{i\in\mathcal{S}}\left(f_{1}\left({\bm{z}}_{i},y_{i},\Theta\right)+f_{2}\left({\bm{z}}_{i},\Theta\right)\right),$ (4)
@@
$f_{1}\left({\bm{z}}_{i},y_{i},\Theta\right)\triangleq\lVert{\bm{z}}_{i}-{\bm{w}}_{y_{i}}\rVert^{2}_{{\mathbf{M}}_{y_{i}}},$ (5)
$f_{2}\left({\bm{z}}_{i},\Theta\right)\triangleq\log\left(\sum_{k\in\{0,1\}}\exp\left(-\lVert{\bm{z}}_{i}-{\bm{w}}_{k}\rVert^{2}_{{\mathbf{M}}_{k}}\right)\right).$
@@
For all $i\in\mathcal{S}$, $f_{1}\left({\bm{z}}_{i},y_{i},\Theta\right)$ is minimized when $\lVert{\mathbf{M}}_{k}\rVert_{F}\to 0$, and $f_{2}\left({\bm{z}}_{i},y_{i},\Theta\right)$ as $\lVert{\mathbf{M}}_{k}\rVert_{F}\to+\infty$.
The function $f_{2}$ is hence responsible for the degenerate solutions obtained while minimizing the cross-entropy loss. We propose to replace $f_{2}$ with a different function $\tilde{f}_{2}$, playing the same role pushing $\lVert{\mathbf{M}}_{k}\rVert_{F}$ towards infinity, but with a less aggressive behaviour. The new loss obtained $f_{1}+\tilde{f}_{2}$ admits a closed-form solution on ${\mathbf{M}}_{k}$, for a fixed ${\bm{w}}_{k}$, which corresponds to the estimator of the empirical covariance matrix’s inverse. The expression of $\tilde{f}_{2}$ is then:
$\tilde{f}_{2}\left(\Theta\right)\triangleq-\sum_{k\in\{0,1\}}\frac{\lvert\mathcal{S}_{k}|\log\det{({\mathbf{M}}_{k})}}{\lvert\mathcal{S}|},$
where $\mathcal{S}_{k}=\{i\in\mathcal{S}|y_{i}=k\}$.